The sum of the ages of Jeff and jay is 42. The sum of the ages of jay and Joseph is 59. The sum of the ages of Jeff and Joseph is 53. Find the ages of Jeff,jay and Joseph .
Answer by addingup(3677)

Jeff: x
Jay : y
Joseph:z
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
x+y = 42
y+z = 59
x+z = 53
~~~~~~~~~~~~~~~~~~~~~~~~~~~~
x+y = 42
-
x+z = 53
--------
y-z = -11
~~~~~~~~~~~~~~~~~~~
y+z = 59
+
y-z = -11
------------
2y = 48 divide both sides by 2
y = 24
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Go back to the first equation and substitute for y:
x+y = 42
x+24 = 42
x = 18
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Now solve for z, take any one of the equations with z in it and solve it, you'll get the same result. Here's proof:
y+z = 59; 24+z = 59; z = 35
x+z = 53; 18+z = 53; z = 35
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
So there you have it:
Jeff: x = 18
Jay : y = 24
Joseph:z = 35
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
Jeff and Jay:18+24 = 42
Jay and Joseph:24+35 = 59
Jeff and Josehp:18+35 = 53
We've got the right answers.
J